OpenStudy (anonymous):

Use the quadratic equation y = x^2 + 3x – 10. Where does the vertex of the parabola lie? A. below the x-axis B. above the x-axis C. on the x-axis D. on the y-axis

OpenStudy (anonymous):

A

OpenStudy (anonymous):

factor that ishh, x^2+3x-10 =(x+5)(x-2) x=-5,2 so you have the points (0,2) and (0,-5) the parabola passes through the x axis twice, and since the x^2 is positive, you know the vertex is below the x axis. the graph curves up like a U your answer is A